TMS Meeting Minutes, December 17, 2020

Attendees: Andrew, Breann, Heejoo, John, Trenton, Matthieu, Olga, Yang (Zoom Meeting)

Meeting Notes:

  • Andrew does not like remote conferences (no opportunities to drink with your pals)
  • John likes remote conferences (on-demand presentations
  • Matthieu working on email laser monitor (expected next week)
    • Planning to inject fake values to test the alarms
  • Will be able to search TMS on DMS to look at the power of the laser monitor
December 10, 2020 Long-Exposure LUCI h-band
  • There were a couple of issues
    • TMS software crashed (2020.12.10 9:37 PM MST)
    • The debug showed nothing
    • Restarted eTalong, ran into "laser device not initialized correctly" error
    • Andrew notes that we should wait two minutes between restarts
    • Heejoo resorted to restarting the PC
    • iMonit showed no accidental shutter opening when the program crashed
    • The system is currently in a crashed state (at present)
    • Andrew working with eTalon engineers to figure out what is going on
    • Fortunately, we have confirmed that we can continuously monitor laser leaking and that the shutter does not open upon eTalon crash
    • The shutters are working well!
  • LUCI Imaging Mode
    • M3/M2 deployed, lasers on, shutter closed
    • No 1550 nm light visible on the detector
    • Unfortunately, imaging mode with the laser on was not conducted
  • Spectroscopic Mode
    • 10 sec exposure
    • Can clearly see the laser line with the shutter open
    • Each pixel is 4 angstroms, wondering if there are +-70 nm lines visible
  • 2018.12.12 UT Images from Olga
    • Two very narrow lines observed
    • Last week, we saw two very broad lines (big difference)
    • It may be that the slit was 1.5", 2"
    • Interesting differences between the images

  • Heejoo suggests we may want to run the lasers continuously for the next month or so to verify effectiveness of shutter
  • Olga mentions that there is a 1.55 micron laser that is used for flexure compensation inside of MODS (internal to MODS)
  • There has been a notable detectable signal from this laser source in MODS images
  • LBC has no detectable signal with the TMS lasers
  • Andrew suggests that we may be able to infer the LUCI and LBTI (1-13 microns) will have equivalent stray light results
  • John asks that Heejoo fix up the presentation to show pixel noise

  • LBC IQ 20201201 Plot
    • Will update with correct temperature measurements
  • John to look into finding a data item for shutter open/close
  • Andrew looking into channel dropping analysis in-depth
    • Not certain as to why the instability is happening
  • Olga notes that donuts are seen under other conditions, as well
    • Notes that temperature changes drive in spherical on LBCr, which could cause donuts
    • Olga wants to check filter focus offset for the red
  • Between 0.5" and 1.5", we have good enough seeing, and that's that
  • Note, TMS can't make spherical aberration
  • LBCb was drifting normally, so John suggests that the drift before 10:30 was caused by TMS

  • John suggests easier way to run this test is not to continuously run FPIA, but to apply a fixed "filter" focus offset (e.g. do the offset for FPIA)
  • Consider post-analyzing with FPIA, don't have to worry about all the little corrections, Z11 and such
  • Just set a 0.8 mm offset + spherical offset to give a real FPIA type image and take a whole bunch of images
  • TMS doesn't mess with the offsets that have already been applied
  • Would be quite valuable for right now, especially in unstable conditions
  • Consider running mirror ventilation if in a super stable condition to push to a less stable condition
Next Week:
  • No meeting next week
  • Updates via email
  • TMS test during break, perhaps
  • Next meeting at the New Year, January 7th
  • If any doubt of the effectiveness of laser shutters, we should just turn off the lasers for now
  • Andrew suggests suspending the ongoing laser on test until we have completed the stray light test on MODS

-- TrentonBrendel - 17 Dec 2020
Topic revision: r1 - 18 Dec 2020, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback